Hi, Im im in the pre-algebra class my teacher gave us a question stating. The are of an right iscolece triangle is 12 squared. What is the approximate length of each of the equal sides she told me this would be solved by an equation.Please help what equation would be needed to solve this problem.
Answer by lwsshak3(11628)

Hi, Im im in the pre-algebra class my teacher gave us a question stating. The area of a right isosceles triangle is 12 squared. What is the approximate length of each of the equal sides she told me this would be solved by an equation.Please help what equation would be needed to solve this problem.
***
As stated, two of the sides of an isosceles triangles are equal.
let x=length of each of these two sides.
draw the triangle so one of the sides(x)is the base(horizontal) and the other side(x)is the height(vertical).
As you know,the formula for the area of a triangle is 1/2*base*height.
The equation, therefore is:
1/2*x*x=12
1/2*x^2=12
multiply each side by lcd:2
x^2=24
x=√24=2√6
Area of given triangle=2√6
note: please learn how to spell isosceles correctly.